Professor Smith took a sample of 196 students from her morning class and a sample of 162 students from her evening class. On Exam 1, all those sampled from the morning class got a 65 while those sampled from the evening class averaged a 72 with a variance of 173.f Professor Smith combines these two samples. Question 2 What is the combined Exam 1 average? Question 3A What is the combined Exam 1median? Question 3B What is the shape of the combined Exam 1 grade distribution? Question 4What is the combined Exam 1standard deviation? HWNotes | HWExcelDirections HW FuelData

PS. Answers with 4 digits
Answer #1

2)combined average=(196*65+162*72)/(196+162)=68.1676

3A)median =65

3B)shape of combined exam I grade distribution: right skewed

4)

here \sumx22 =(n-1)*s2+(\sumx2))2/n =(162-1)*173+(72*162)2/162=867661

hence combined sum =\sumx =65*196+72*162=24404

also sum of square \sumx2 =\sumx12 +\sumx22 =196*652+27885=855985

hence std deviation =sqrt( (\sumx2-( \sumx)2/n)/(n-1))=9.4970